OpenStudy (anonymous):

HELP PLEASE!!! An elegant men's shop is having a post Christmas sale. All shirts are reduced to one low price, and all ties are reduced to one lower price. A customer purchased 3 ties and 7 shirts, paying $570. Another customer selects 5 ties and 3 shirts and pays $430. what is the sale price of each tie and of each shirt? CAN SOMEONE PLEASE HELP ME BY SHOWING ME STEP BY STEP HOW TO WORK THIS PROBLEM!!!!

OpenStudy (anonymous):

Okay first I would create an equation out of this word problem here's mine 3x+7y=570 and 5x+3y=430, where x=ties and y=shirts. Now what you have to do is find a value for x and y that work in each equation.

OpenStudy (anonymous):

Hopefully Wired can continue this

OpenStudy (anonymous):

Where ties are X and shirts are Y: Three ties and 7 shirts are 570. This can be written as: 3X + 7Y = 570 5 ties and 3 shirts is 430. This can be written as: 5X + 3Y = 430 Solve the first equation for X. Then substitute that value (which will be a formula) into the second equation. That will get you Y. Take that value (which will be a number) and substitute it for Y into either equation and you have your answer.

OpenStudy (anonymous):

OK......

OpenStudy (anonymous):

Understand? or need more help

OpenStudy (anonymous):

can you show me what he said:Solve the first equation for X. Then substitute that value (which will be a formula) into the second equation. That will get you Y. Take that value (which will be a number) and substitute it for Y into either equation and you have your answer.

OpenStudy (anonymous):

im horrible in math :(

OpenStudy (anonymous):

When someone says: "solve for a variable", they mean you need to manipulate the equation so that the variable is by itself on one side of =. Have you done this yet in class?

OpenStudy (anonymous):

no i haven't

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

Not sure why they assigned this to you then!

OpenStudy (anonymous):

i know i just want to see how you do this problem so i can have a guide and do the other ones by myself

OpenStudy (anonymous):

Agreed

OpenStudy (anonymous):

I see well give me a sec its hard to explain this on a computer

OpenStudy (anonymous):

ok i can wait. thank you!

OpenStudy (anonymous):

Think of an equation as the type of scale where both sides have to balance. To modify one side you have to do the same thing to the other side. 3X + 7Y = 570 Subtract 3X from both sides: 3X+7Y - 7Y = 570 - 7Y Simplify the left side: 3X + (7Y - 7Y) = 570 - 7Y Since anything minus itself is 0, 7Y - 7Y = 0. Replace (7Y - 7Y) with 0 in the equation and you get: 3X + 0 = 570 - 7Y Since (anything) + 0 = (anything): 3X = 570 - 7Y Again, you have to do the same thing to both sides, so divide both sides by 3 (same thing as multiplying both sides by 1/3): (3X)/3 = (570 - 7Y)/3 Simplify the equation gets you: X = (570 - 7Y)/3

OpenStudy (anonymous):

Okay what wired meant was once you have the equations 3x+7y=570 and 5x+3y=430, you should solve for one of the variables that way you can substitute the equation into the other (this is the only way to solve these equations other than guess and check). so I will solve for y in 5x+3y=430 It will go like this 5x+3y=430 subtract 5x from both sides 3y=430-5x then divide 3 from both sides y=(430-5x)/3 don, now substitute 3x+7(430-5x)/3=570 simplify 3x+(3010-35x)/3=570 3(3x+(3010-35x)/3)=560(3) 9x+3010-35x=1680 -27x+3010=1680 subtract 3010 from both sides -27x=-1330 now divide both sides by -27 x=-1330/-27 which simplifies to 49.259259 now in money this should be $49.26 for x or ties now substitute our answer into wither equation 3(49.26)+7y=570 and solve (I hope you know how to do this) the answer should be y=60.317 which in money should be $60.32 Your answers are 60.32 for shirts and 49.26 for ties Now the problem with computers is I can't properly show you why I did here what I did also I'm sure I might have made a arithmatic error but the above process is correct. To others if you mind adding to my answer it would be much appreciated and would help colombianita19 Thanks.

OpenStudy (anonymous):

Great! thanks! :)

OpenStudy (anonymous):

I'm glad I help, lol, I felt like I was making the problem worse

OpenStudy (anonymous):

nah! thanks! :)